OpenStudy (anonymous):

pair of functions, f(x)=x^2+4 and g(x)=x+5 find the following - (f little tiny o g)(x) so its looks like(fog)(x), then (fog)(2), and then (gof)(x) last one (gof)(2)

OpenStudy (anonymous):

(fog)(x)=(x+5)^2+4 and that simplifies to x^2 + 10x + 29

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

f(x)=x^2+4 and g(x)=x+5 find the following (gof)(x)